OpenStudy (anonymous):

CHECK MY ANSWER What is the product written in scientific notation? (1.4 x 10^3)(2.5 x 10^-6) A. 3.9 x 10^-3 B. 3.9 x 10^-18 C. 3.5 x 10^-3 D. 3.5 x 10^-2 <<<<

OpenStudy (anonymous):

Nope You have the mantissa correct but the wrong exponent.

OpenStudy (welshfella):

multiply the first parts first - that is 1.4 * 2.5 then use the laws of indices to calculate the powers 10^a * 10^b = 10^(a + b)

OpenStudy (tkhunny):

Please find a better way to format your questions. We don't need all that white space. Show how you got your answer!

OpenStudy (anonymous):

i would say yes your correct

TheSmartOne (thesmartone):

Also copy and pasting fail, you lost all the powers to.

OpenStudy (welshfella):

10^3 * 10*-6 = 10^(3 - 6)

OpenStudy (welshfella):

* that should be 10^-6

OpenStudy (welshfella):

which option do you think is correct?